# Effect of dark matter on the shadow of a distorted and deformed compact object

### v1

@@ III Shadows and light rings
The black hole shadow for a far-away observer is defined as the set of directions in the sky, propagated backwards in time, that cross the event horizon. Its boundary consists of those rays that asymptotically approach the spherical photon orbits of the underlying spacetime, i.e., the photon spheres (PS) or critical curves, with the same specific energy and angular momentum that are conserved quantities. In fact, the shape and size of the shadow depend on the geometry of the background spacetime, which requires the analysis of the light-ray or null geodesics of a test particle.
Figure 1: Luminosity profiles for the three models of accretion disks considered, i.e., the ISCO model (red solid), the LR model (blue dashed), and the EH model (purple dotted), for $\alpha=0$ and $\beta=0$.
III.1 Null geodesics and impact parameter
@@
$\mathcal{L}=\frac{1}{2}\left[-g_{tt}\dot{t}^{2}+g_{rr}\dot{r}^{2}+g_{\phi\phi}\dot{\phi}^{2}\right]$ (4)
@@
$E=\frac{\partial\mathcal{L}}{\partial\dot{t}}=g_{tt}\dot{t},\quad\quad L=\frac{\partial\mathcal{L}}{\partial\dot{\phi}}=g_{\phi\phi}\dot{\phi}.$ (5)
@@
For a given photon propagating in the background spacetime, if $\dot{r}$ vanishes at a radius larger than the critical curve, this implies that the photon reaches a radius of closest approach and subsequently propagates back to infinity. If this situation is never verified, then the photon is captured by the central object. Introducing the impact parameter $b=\frac{L}{E}$ and rewritting Eq. (6), one can obtain the so-called critical impact parameter $b_{\rm crit}$, which discriminates between captured and uncaptured photons. This can be done by solving the equations $\dot{r}=0$ and $\ddot{r}=0$ for circular orbits, which leads to
@@
The signs ($\pm$) represent ingoing (--) and outgoing ($+$) null geodesics. In what follows, we perform a numerical integration of Eq. (9) using a Mathematica-based ray-tracing code that has been previously used in several publications to simulate the observational properties of compact objects ([25, 26, 27, 52, 53, 54, 55, 56, 57, 28, 58]). Note that a given ingoing geodesic transitions into an outgoing geodesic at the radius of the closest approach, i.e., when $d\phi/dr$ diverges, at which point one should invert the sign of Eq. (9) to proceed with the integration. By employing the spherically symmetric setup, considering relatively small quadrupoles and restricting the analysis to the equatorial plane $\theta=\frac{\pi}{2}$, the study of circular geodesics i.e. $\dot{r}=0$ and $\ddot{r}=0$, the radius of these circular geodesics is related to $\alpha$ and $\beta$ through the following relation[37]
@@
where in the limit $\beta=0$ one recovers the radius of the PS for the $\rm q$-metric $r_{\rm LR}=M(3+2\alpha)$, and in the limit $\alpha=\beta=0$ one recovers the Schwarzschild radius of the PS $r_{\rm LR}=3M$. In addition, one can have one bound photon sphere in the equatorial plane for any given valid $\beta<-\frac{1}{2}$ with
@@
This is not the case neither in the Schwarzschild spacetime, nor in the $\rm q$-metric. In fact, this arises due to the existence of quadrupoles related to the external source and indicating that, for certain negative values of $\beta$, a bound photon orbit exists in the equatorial plane. It is necessary to exclude these orbits in order to conduct a study on shadows.
III.2 Accretion flow models and intensity profiles
@@
We assume that the accretion disk is located in the static rest frame and emits isotropically. More importantly, the whole setup should be close enough to the horizon since photons emitted from this vicinity with a small enough impact parameter are absorbed by the central compact object and cast a shadow. In fact, the inner part of the accretion disc flow has the largest density, therefore most of the radiation comes from this region with the highest temperature. If the disk has no magnetic field with sub-Eddington luminosities then, with a high accuracy, the place of the innermost stable circular orbit (ISCO) is considered as the inner edge of the thin disk [29] where most of the luminosity comes from. Therefore, it can be used to infer possible luminosity and intensity profiles for different accretion disk models. However, in what follows, we consider emission from an optically thin and geometrically thin region near the compact object, with three different stationary accretion disk models. In addition, the observer stands on the vertical axis $\theta=0$. Under those considerations, the intensity models are as follows:
$\beta$ $\alpha$ $-0.3$ 00 $0.3$
$-10^{-5}$
$4.29995M$
$2.40003M$
$5.98462M$
$3.00012M$
$7.59323M$
$3.60030M$
@@
  3. 3.
    The EH model: In this model, the emissivity extends to the event horizon, and is not truncated at any larger radius. This model is motivated by the expectation that the region $r<r_{\rm LR}$ is populated by matter falling into the central compact object all the way down to the Event Horizon (EH), which for the Schwarzschild spacetime ($\alpha=\beta=0$) stands at $r_{\rm EH}=2M$. We thus consider an additional intensity model for which the luminosity increases monotonically from $r=\infty$ and peaks at $r_{\rm EH}$.
@@
Figure 2: Shadow images for the ISCO disk model with different combinations of $\alpha$ and $\beta$, namely $\alpha=-0.3$ (top row), $\alpha=0$ (middle row), and $\alpha=+0.3$ (bottom row); and $\beta=-10^{-5}$ (left column), $\beta=0$ (middle column), and $\beta=10^{-5}$ (right) column).
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Note that for $\alpha\neq 0$ the spacetime does not have an event horizon but a naked singularity at $r=2M$. However, for the sake of notation simplicity, we stick to the symbol $r_{\rm EH}=2M$ and pictorial representation for these solutions. To model the three intensity profiles described above, we recur to an emission profile published recently in [18] and shown to be in close agreement with observational predictions and general relativistic magneto-hydrodynamics (GRMHD) simulations. The intensity profile $I\left(r\right)$ of this model is derived from Johnson’s SU distribution, which is a free function with a smooth profile, mostly concentrated within a few Schwarzschild radii of the central object, and increases as we approach the compact object. This emission profile is given by
@@
where $\gamma$, $\mu$, and $\sigma$ are free parameters that shape the behavior of the function $I\left(r\right)$. The parameter $\gamma$ controls the rate of increase of the luminosity from infinity down to the peak; the parameter $\mu$ induces a translation of the profile and controls the radius at which the luminosity peaks; and the parameter $\sigma$ controls the dilation of the profile. The values of the parameters $\gamma$, $\mu$ and $\sigma$ for each of the three disk models mentioned above, i.e., the ISCO, LR and EH disk models, are specified in Table 2, where $r_{LR}$ and $r_{ISCO}$ are provided in Table 1, and the respective intensity profiles are plotted in Figure 1. In this Figure, different values of $\alpha$ and $\beta$ are not plotted separately as they correspond solely to a translation of the ISCO and LR models in the horizontal axis. However, we adapted these models in our set-up considering the place of ISCO and light ray is changing with $\alpha$ and $\beta$ [37].
Note that the intensity profiles plotted in Figure 1 correspond to the intensity profiles in the reference frame of the accretion disk. The frequency in the reference frame of the observer and the emitter, are denoted by $\nu_{o}$ and $\nu_{e}$, respectively. These two frequencies are related via the gravitational redshift as $\nu_{0}=\sqrt{g_{tt}}\nu_{e}$, then the associated specific intensity $i\left(\nu\right)$ scales as
@@
where $n=\frac{\theta}{2\pi}$, and $r_{n}(b)$ is the demagnification factor $\frac{dr}{db}$ which represents the intersection radius with the equatorial plane for the nth times on its backward journey, where $b$ is the impact parameter. The continuous version of Eq. (14) is also taken into consideration in our Mathematica-based code to produce the intensity profiles in the reference frame of the observer.
Figure 4: Shadow images for the LR disk model with different combinations of $\alpha$ and $\beta$, namely $\alpha=-0.3$ (top row), $\alpha=0$ (middle row), and $\alpha=+0.3$ (bottom row); and $\beta=-10^{-5}$ (left column), $\beta=0$ (middle column), and $\beta=10^{-5}$ (right) column).
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
III.3 Shadows and observed intensity profiles
@@
To analyze the effects of the parameters $\alpha$ and $\beta$ in the observational properties of the background spacetime given in Eq.(II) in the regime where the spherically symmetric approximation is valid, we have run the ray-tracing code for several combinations of setups. In particular, we have considered, $\alpha=\{-0.3;0;0.3\}$, and $\beta=\{-10^{-5};0;10^{-5}\}$, for the three accretion disk luminosity profiles given in Table 2. The accretion disk is placed at the equatorial plane, i.e., at $\theta=\pi/2$, while the observer is placed at the vertical axis $\theta=0$ and at a radius of $r=600M$, far enough from the central object such that the light rays are approximately parallel when they reach the observer. Note that this solution is valid locally due to the existence of external fields encoded in $\beta$. In the equatorial plane validity region for $\beta<0$ is wider than for $\beta>0$, and an increase in $\beta$ leads to a decrease in the size of this region. We also chose relatively small values of $\beta$ with the goal of having a wider validity region allowing for the comparison of the results among different cases. The shadow images and the influence of $\alpha$ and $\beta$ in the observed intensity profiles are given in Figures 2 and 3 for the ISCO model, Figures 4 and 5 for the LR model, and Figure 6 and 7 for the EH model. As for photon rings, their size and location can be quite different depending on the features of the parameters. In addition, the corresponding luminosity is also strongly related to the background geometry and the properties of the disk.
@@
The first important feature to note is that the observed intensity profiles feature multiple peaks of intensity, in contrast to the emitted intensity profiles who feature a single peak of intensity, see Figure 1. These multiple peaks of intensity at the observer are caused by photons that reach the observer after contouring the central object with a different number of half-orbits. In this peak structure, we can identify three main components:
@@
    the direct component: corresponding to the wider and dominant component, it is composed of the photons emitted directly from the accretion disk at $\theta=\pi/2$, i.e., the equatorial plane, to the observer at $\theta=0$. These photons have traveled a total of a quarter orbit around the central object, $\Delta\theta=\theta_{e}-\theta_{o}=\pi/2$, and is defined as the primary image;
@@
    the lensed component: corresponding to a narrower component placed either inside (for the ISCO model) or outside (for the LR and EH models) the inner edge of the direct component, is composed of the photons that have been emitted by the accretion disk in the direction opposite to the observer, and then have been lensed around the central object once before reaching the observer. These photons have traveled an extra half-orbit than the photons of the direct component, $\Delta\theta=3\pi/2$, and is defined as the secondary image;
@@
    the photon ring component: corresponding to the narrowest of the three components, it is composed of the photons that have been emitted by the accretion disk and have orbited at least once around the central object, close to the LR, before reaching the observer. This component presents an infinite sub-structure of peaks corresponding to the photons that have orbited a total of $\Delta\theta=5\pi/2+n\pi$, for $n\geq 0$ an integer, which leads to additional brightness once again. This infinite sub-structure of peaks converges to critical curve $b=b_{c}$ which becomes exponentially closer to each other with an increase in the number of orbits.
@@
For the ISCO model, these three components are all discretized in the observed emission profiles, see e.g. the central image in Figure 2 for $\alpha=0$ and $\beta=0$. However, for the LR model, the direct and lensed components appear superimposed, and for the EH model, the photon ring component is also superimposed with the direct emission.
@@
In all three models considered for the accretion disk, one verifies that the parameter $\alpha$ affects the results more prominently than the parameter $\beta$. Indeed, since a variation in $\alpha$ induces a non-negligible translation of the radii of the ISCO and the LR of the background spacetime, one observes that the size of the shadow is strongly affected by this parameter, with the shadow radius increasing for $\alpha>0$ and decreasing for $\alpha<0$. Furthermore, one also verifies that the maximum amplitude of the observed intensity profiles for the direct emission is inversely proportional to the value of $\alpha$. This behavior is expected as, for negative values of both quadrupole parameters, the radius of the ISCO is smaller than for the positive values [37]; thus the radiation is being emitted closer to the central object. Indeed, for all of the accretion disk models, one verifies that the rate of change of the intensity profile as a function of the radial coordinate is more abrupt for larger values of $\alpha$ which is also expected. As a consequence, the lensed and photon ring components of the intensity profiles for negative values of alpha are wider, and thus more visible in the shadow images. On the other hand, the parameter $\beta$ presents a minor effect on the observational properties of this setting, inducing smaller changes in the observed intensity profiles in comparison with the effect of the parameter $\alpha$. Our results show that the observed intensity profiles are slightly closer to the central object for negative values of $\beta$, while positive values of beta tend to bring them farther. The effects are particularly clear for the ISCO model, for which the direct and lensed components of the observed intensity profile are separated, but it is also visible (although less contrasting) in the LR and EH models, for which the direct and lensed components of the observed intensity profile are superimposed.
@@
Figure 8 shows how the null geodesic congruences are affected by parameters $\alpha$ and $\beta$. The left panel shows the effect of $\alpha$ for a wide range of impact parameters. Most trajectories cross the equatorial plane once, but photons that propagate close to the critical curve on the image plane wrap around the central object and cross the plane twice, three times or more, with photons appearing exactly on the critical curve describing trajectories that asymptotically approach the unstable bound orbits composing the photon sphere. The size of the shadow is defined by the range of impact parameters corresponding to the trajectories that do not cross the equatorial plane before intersecting the black disk.Thus, one can see that the size of the shadow produced by an oblate compact object is larger than the shadow produced by the Schwarzschild blackhole, and the reverse is true for a prolate compact object. In the right panel, the impact of the parameter $\beta$ is shown. It is evident that the effect of the parameter $\beta$ in this setting and in the equatorial plane is much smaller in comparison with the effects of the parameter $\alpha$ as anticipated by the analysis of the intensity profiles and shadow images.
Figure 6: Shadow images for the EH disk model with different combinations of $\alpha$ and $\beta$, namely $\alpha=-0.3$ (top row), $\alpha=0$ (middle row), and $\alpha=+0.3$ (bottom row); and $\beta=-10^{-5}$ (left column), $\beta=0$ (middle column), and $\beta=10^{-5}$ (right) column).
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 7: Intensity $I(r)$ as a function of the radius $r$ for the EH disk model for a varying $\alpha$ with constant $\beta$ (top row) and for a varying $\beta$ with constant $\alpha$ (bottom row)
@@
Figure 8: Photon trajectories that reach a distant observer in the far right $z\to\infty$. The black disk at $x=1$ represents the event horizon for $\alpha=0$ or the naked singularity for $\alpha\neq 0$. Photon trajectories are colored according to the different parameters of the metric. Left panel: $\beta=0$ and prolate object $\alpha=-0.3$ (blue lines), Schwarzschild solution $\alpha=0$ (black lines) and oblate object $\alpha=0.3$ (red lines). Right panel: $\alpha=0$ and $\beta=-10^{-5}$ (blue lines), Schwarzschild solution $\beta=0$ (black lines), and $\beta=10^{-5}$ (red lines).
IV Dark matter model

### v2

@@ 3 Shadows and light rings
The black hole shadow for a far-away observer is defined as the set of directions in the sky, propagated backwards in time, that cross the event horizon. Its boundary consists of those rays that asymptotically approach the surfaces of the underlying spacetime on which photons undergo unstable closed orbits, i.e., the photon shell (PS) or critical curves, with the same specific energy and angular momentum that are conserved quantities. In fact, the shape and size of the shadow depend on the geometry of the background spacetime, which requires the analysis of the light-ray or null geodesics of a test particle.
3.1 Null geodesics and impact parameter
@@
$\mathcal{L}=\frac{1}{2}\left[g_{tt}\dot{t}^{2}-g_{rr}\dot{r}^{2}-g_{\phi\phi}\dot{\phi}^{2}\right]$ (4)
@@
$E=\frac{\partial\mathcal{L}}{\partial\dot{t}}=g_{tt}\dot{t},\quad\quad L=\frac{\partial\mathcal{L}}{\partial\dot{\phi}}=-g_{\phi\phi}\dot{\phi}.$ (5)
@@
For a given photon propagating in the background spacetime, if $\dot{r}$ vanishes at a radius larger than the unstable photon orbit, this implies that the photon reaches a radius of closest approach and subsequently propagates back to infinity. If this situation is never verified, then the photon is captured by the central object. Introducing the impact parameter $b=\frac{L}{E}$ and rewriting Equation (6), one can obtain the so-called critical impact parameter $b_{\rm crit}$, which discriminates between captured and uncaptured photons. This can be done by solving the equations $\dot{r}=0$ and $\ddot{r}=0$ for circular orbits, which leads to
@@
The signs ($\pm$) represent ingoing (--) and outgoing ($+$) null geodesics. In what follows, we perform a numerical integration of Eq. (9) using a Mathematica-based ray-tracing code that has been previously used in several publications to simulate the observational properties of compact objects ((Rosa & Rubiera-Garcia, 2022; Rosa et al., 2023; Rosa, 2023; Olmo et al., 2023; Guerrero et al., 2022b, a; Olmo et al., 2022; Guerrero et al., 2021; Asuküla et al., 2024; Rosa et al., 2024; Macedo et al., 2024)). Note that a given ingoing geodesic transitions into an outgoing geodesic at the radius of the closest approach, i.e., when $d\phi/dr$ diverges, at which point one should invert the sign of Eq. (9) to proceed with the integration. Note that, contrary to what happens in spherically symmetric spacetimes, the right-hand side of Eq. (9) depends explicitly on complicated functions of $\theta$, which is a major obstacle in the numerical resolution of the equation. To simplify this procedure and improve the computational time, we employ the spherically symmetric approximation, i.e., we consider small values of the quadrupole parameters and we restrict the analysis to the equatorial plane $\theta=\frac{\pi}{2}$. The errors incurred by taking this approximation are always smaller than $1\%$ for the situations analysed in this work (see Appendix A for more details on the validity of this approximation). The study of circular geodesics i.e. $\dot{r}=0$ and $\ddot{r}=0$, the radius of these circular geodesics is related to $\alpha$ and $\beta$ through the following relation(Faraji, 2022)
@@
where in the limit $\beta=0$ one recovers the radius of the PS for the $\rm q$-metric $r_{\rm LR}=M(3+2\alpha)$, and in the limit $\alpha=\beta=0$ one recovers the Schwarzschild radius of the PS $r_{\rm LR}=3M$. In addition, one can have one bound unstable photon orbit in the equatorial plane for any given valid $\beta<-\frac{1}{2}$ with
@@
This is not the case either in the Schwarzschild spacetime, nor in the $\rm q$-metric. In fact, this arises due to the existence of a quadrupole related to the external source and indicating that, for certain negative values of $\beta$, an unstable bound photon orbit exists in the equatorial plane. It is necessary to exclude these orbits in order to conduct a study on shadows.
3.2 Accretion flow models and intensity profiles
@@
We assume that the accretion disk is located in the static rest frame and emits isotropically. More importantly, the whole setup should be close enough to the horizon since photons emitted from this vicinity with a small enough impact parameter are absorbed by the central compact object and cast a shadow. In fact, the inner part of the accretion disc flow has the largest density, therefore, most of the radiation comes from this region with the highest temperature. If the disk has no magnetic field with sub-Eddington luminosities, then, with high accuracy, the place of the innermost stable circular orbit (ISCO) is considered as the inner edge of the thin disk (Afshordi & Paczyński, 2003) where most of the luminosity comes from. Therefore, it can be used to infer possible luminosity and intensity profiles for different accretion disk models. However, in what follows, we consider emission from an optically thin and geometrically thin region near the compact object, with three different stationary accretion disk models. In addition, the observer stands on the vertical axis $\theta=0$. Under those considerations, the intensity models are as follows:
$\beta$ $\alpha$ $-0.01$ 00 $0.01$
$-10^{-6}$
$5.94346M$
$2.98001M$
$5.99844M$
$3.00001M$
$6.05336M$
$3.02001M$
@@
  3. III.
    The EH model: In this model, the emissivity extends to the event horizon and is not truncated at any larger radius. This model is motivated by the expectation that the region $r<r_{\rm LR}$ is populated by matter falling into the central compact object all the way down to the Event Horizon (EH), which for the Schwarzschild spacetime ($\alpha=\beta=0$) stands at $r_{\rm EH}=2M$. We thus consider an additional intensity model for which the luminosity increases monotonically from $r=\infty$ and peaks at $r_{\rm EH}$.
@@
Note that for $\alpha\neq 0$ the spacetime does not have an event horizon but a naked singularity at $r=2M$. However, for the sake of notation simplicity, we stick to the symbol $r_{\rm EH}=2M$ and pictorial representation for these solutions. To model the three intensity profiles described above, we recur to an emission profile published recently in (Gralla et al., 2020) and shown to be in close agreement with observational predictions and general relativistic magneto-hydrodynamics (GRMHD) simulations. The intensity profile $I\left(r\right)$ of this model is derived from Johnson’s SU distribution, which is a free function with a smooth profile, mostly concentrated within a few Schwarzschild radii of the central object, and increases as we approach the compact object. This emission profile is given by
@@
where $\gamma$, $\mu$, and $\sigma$ are free parameters that shape the behavior of the function $I\left(r\right)$. The parameter $\gamma$ controls the rate of increase of the luminosity from infinity down to the peak; the parameter $\mu$ induces a translation of the profile and controls the radius at which the luminosity peaks; and the parameter $\sigma$ controls the dilation of the profile. The values of the parameters $\gamma$, $\mu$ and $\sigma$ for each of the three disk models mentioned above, i.e., the ISCO, LR and EH disk models, are specified in Table 2, where $r_{LR}$ and $r_{ISCO}$ are provided in Table 1, and the respective intensity profiles are plotted in Figure 2. In this Figure, different values of $\alpha$ and $\beta$ are not plotted separately as they correspond solely to a translation of the ISCO and LR models in the horizontal axis. However, we adapted these models in our set-up considering the place of ISCO and light ray is changing with $\alpha$ and $\beta$ (Faraji, 2022).
Figure 2: Luminosity profiles for the three models of accretion disks considered: the ISCO model (red solid), the LR model (blue dashed), and the EH model (purple dotted), for $\alpha=0$ and $\beta=0$.
Note that the intensity profiles plotted in Figure 2 correspond to the intensity profiles in the reference frame of the accretion disk. The frequencies in the reference frame of the observer and the emitter, are denoted by $\nu_{o}$ and $\nu_{e}$, respectively. These two frequencies are related via the gravitational redshift as $\nu_{0}=\sqrt{g_{tt}}\nu_{e}$, then the associated specific intensity $i\left(\nu\right)$ scales as
@@
where $n=\frac{\theta}{2\pi}$, and $r_{n}(b)$ is the demagnification factor $\frac{dr}{db}$ which represents the intersection radius with the equatorial plane for the nth time on its backwards journey, where $b$ is the impact parameter. The continuous version of Eq. (14) is also taken into consideration in our Mathematica-based code to produce the intensity profiles in the reference frame of the observer.
3.3 Shadows and observed intensity profiles
@@
To analyse the effects of the parameters $\alpha$ and $\beta$ in the observational properties of the background spacetime given in Eq.(2) in the regime where the spherically symmetric approximation is valid, we have run the ray-tracing code for several combinations of setups. In particular, we have considered, $\alpha=\{-0.01;0;0.01\}$, and $\beta=\{-10^{-5};0;10^{-5}\}$, for the three accretion disk luminosity profiles given in Table 2. The accretion disk is placed at the equatorial plane, i.e., at $\theta=\pi/2$, while the observer is placed at the vertical axis $\theta=0$ and at a radius of $r=600M$. Note that this approximation is valid only locally due to the alterations in the metric components caused by the parameter $\beta$. These alterations cause the spacetime to lose its asymptotic flatness, and thus the distance from the central object to the observer can not be made arbitrarily large. The value $r=600M$ was chosen in such a way that the effects of losing the asymptotic flatness are still negligible, while maintaining a distance large enough for the light rays to be approximately parallel when they reach the observer. In the equatorial plane validity region for $\beta<0$ is wider than for $\beta>0$, and an increase in $\beta$ leads to a decrease in the size of this region. We also chose relatively small values of $\beta$ with the goal of having a wider validity region allowing for the comparison of the results among different cases. Our results indicate that the effects of the parameter $\beta$ are negligible on the observed shadow image and luminosity profiles. Thus, in this section, we restrict our analysis to the effects of the parameter $\alpha$. For a complete set of results including the effects of the parameter $\beta$, we refer the reader to Appendix B. The shadow images for the three disk models are presented in Figure 3, whereas the observed intensity profiles for the three disk models are presented in Figure 4. As for photon rings, their size and location can be quite different depending on the features of the parameters. In addition, the corresponding luminosity is also strongly related to the background geometry and the properties of the disk.
Figure 3: Shadow images for the ISCO disk model (top row), the LR disk model (middle row), and the EH disk model (bottom row), for $\alpha=-0.01$ (left column), $\alpha=0$ (middle column), and $\alpha=+0.01$ (right column).
Figure 4: Observed intensity $I(r)$ as a function of the radius $r$ for a varying $\alpha$ with constant $\beta=0$, and for the ISCO model (left panel), LR model (middle panel), and the EH model (right panel).
@@
The first important feature to note is that the observed intensity profiles feature multiple peaks of intensity, in contrast to the emitted intensity profiles which feature a single peak of intensity, see Figure 2. These multiple peaks of intensity at the observer are caused by photons that reach the observer after circling around the central object with a different number of half-orbits. In this peak structure, we can identify three main components:
@@
    the direct component: corresponding to the wider and dominant component, it is composed of the photons emitted directly from the accretion disk at $\theta=\pi/2$, i.e., the equatorial plane, to the observer at $\theta=0$. These photons have travelled a total of a quarter orbit around the central object, $\Delta\theta=\theta_{e}-\theta_{o}=\pi/2$, and are defined as the primary image;
@@
    the lensed component: corresponding to a narrower component placed either inside (for the ISCO model) or outside (for the LR and EH models) the inner edge of the direct component, is composed of the photons that have been emitted by the accretion disk in the direction opposite to the observer, and then have been lensed around the central object once before reaching the observer. These photons have travelled an extra half-orbit than the photons of the direct component, $\Delta\theta=3\pi/2$, and are defined as the secondary image;
@@
    the photon ring component: corresponding to the narrowest of the three components, it is composed of the photons that have been emitted by the accretion disk and have orbited at least once around the central object, close to the LR, before reaching the observer  [^2] We note that, due to the limited resolution of our images and the fact that the photon ring component is the thinnest of the three observable components, the photon ring component should be used solely to determine the position of the critical curve in the observed image, and not to trace conclusions about the observed intensity.. This component presents an infinite sub-structure of peaks corresponding to the photons that have orbited a total of $\Delta\theta=5\pi/2+n\pi$, for $n\geq 0$ an integer, which leads to additional brightness once again. This infinite sub-structure of peaks converges to critical curve $b=b_{c}$ which becomes exponentially closer to each other with an increase in the number of orbits.
@@
For the ISCO model, these three components are all distinguishable in the observed emission profiles, see e.g., the top row in Figure 3 and the left panel in Figure 4. However, for the LR model, the direct and lensed components appear superimposed, and for the EH model, the photon ring component is also superimposed with the direct emission.
@@
In all three models considered for the accretion disk, one verifies that the parameter $\alpha$ affects the results more prominently than the parameter $\beta$. Indeed, since a variation in $\alpha$ induces a non-negligible translation of the radii of the ISCO and the LR of the background spacetime, one observes that the size of the shadow is strongly affected by this parameter, with the shadow radius increasing for $\alpha>0$ and decreasing for $\alpha<0$. Furthermore, one also verifies that the maximum amplitude of the observed intensity profiles for the direct emission is anti-correlated to the value of $\alpha$. This behaviour is expected as, for negative values of both quadrupole parameters, the radius of the ISCO is smaller than for the positive values (Faraji, 2022); thus, the radiation is being emitted closer to the central object. Indeed, for all of the accretion disk models, one verifies that the rate of change of the intensity profile as a function of the radial coordinate is more abrupt for larger values of $\alpha$, which is also expected. As a consequence, the lensed and photon ring components of the intensity profiles for negative values of alpha are wider, and thus more visible in the shadow images. On the other hand, the parameter $\beta$ presents a minor effect on the observational properties of this setting, inducing smaller changes in the observed intensity profiles in comparison with the effect of the parameter $\alpha$, see Appendix B. Our results show that the observed intensity profiles are slightly closer to the central object for negative values of $\beta$, while positive values of beta tend to bring them farther. The effects are particularly clear for the ISCO model, for which the direct and lensed components of the observed intensity profile are separated, but it is also visible (although less contrasting) in the LR and EH models, for which the direct and lensed components of the observed intensity profile are superimposed.
@@
Figure 5 shows how the null geodesic congruences are affected by parameters $\alpha$ and $\beta$. Although all calculations and plots extend up to an observer distance of $600M$, in this Figure we only display results out to $100M$ for convenience. The left panel shows the effect of $\alpha$ for a wide range of impact parameters. Most trajectories cross the equatorial plane once, but photons that propagate close to the unstable photon orbit on the image plane wrap around the central object and cross the plane twice, three times or more, with photons appearing exactly on the critical curve describing trajectories that asymptotically approach the unstable bound orbits composing the photon shell. The size of the shadow is defined by the range of impact parameters corresponding to the trajectories that do not cross the equatorial plane before intersecting the black disk. Thus, one can see that the size of the shadow produced by an oblate compact object is larger than the shadow produced by the Schwarzschild blackhole, and the reverse is true for a prolate compact object. In the right panel, the impact of the parameter $\beta$ is shown. It is evident that the effect of the parameter $\beta$ in this setting and in the equatorial plane is much smaller in comparison with the effects of the parameter $\alpha$ as anticipated by the analysis of the intensity profiles and shadow images.
@@
Figure 5: Photon trajectories that reach a distant observer in the far right $z=100M$. The black disk at $x=1$ represents the event horizon for $\alpha=0$ or the naked singularity for $\alpha\neq 0$. Photon trajectories are colored according to the different parameters of the metric. Left panel: $\beta=0$ and prolate object $\alpha=-0.01$ (blue lines), Schwarzschild solution $\alpha=0$ (black lines) and oblate object $\alpha=0.01$ (red lines). Right panel: $\alpha=0$ and $\beta=-10^{-6}$ (blue lines), Schwarzschild solution $\beta=0$ (black lines), and $\beta=10^{-6}$ (red lines). Note that each set of three geodesics is emitted with the same impact parameter at $z=100M$ but, due to the fact that the spacetime is not asymptotically flat, they reach the vicinity of the central object visible on the plot with slight deviations from the asymptotically flat $\alpha=0$ or $\beta=0$ cases.
4 Dark matter model